OpenStudy (anonymous):

Solve g(f(x)) when f(x)=x^2+5 and g(x)=squareroot of 9-x and can you please help me with the problem i know how to do it i just cant find the answer!

OpenStudy (hunus):

\[g(f(x))\] \[f(x)=x^2+5\] \[g(x)=\sqrt{9-x}\] \[g(x^2+5)=\sqrt{9-(x^2+5)}=\sqrt{4-x^2}\] or \[\sqrt{2-x}\sqrt{2+x}\]